OpenStudy (anonymous):

determine the volume of the solid of revolution generated by revolving the region bounded by y=x^3-x^5, y=0, x=0 and x=1 about the line x=3

OpenStudy (tkhunny):

Directly or by Pappus?

OpenStudy (anonymous):

directly

OpenStudy (tkhunny):

\(\pi\cdot\int\limits_0^1 Distance\;dx\) Distance = 3 - y